Tickets for the homecoming dance cost $20 for a single ticket or $35 for a couple. Ticket sales totaled $2280, and 128 people attended. How many tickets of each type were sold? Be sure to define your variables first.

1
Expert's answer
2021-11-30T12:33:29-0500

x - the number of single ticket

y - he number of couple ticket

"x + 2y=128 \to x = 128-2y"

"20x+35y=2280"

"20(128-2y)+35y=2280"

"2560-40y+35y=2280"

"5y=280 \to y = 56 \space \space and \space \space x = 128-2y=16"

Answer

56 - the number of single ticket

16 - the number of couple ticket